OpenStudy (ashley1nonly):

For the reaction: N2(g) + 2 O2(g) ⇌ 2 NO2(g), Kc = 8.3 × 10-10 at 25°C. What is the concentration of N2 gas at equilibrium when the concentration of NO2 is five times the concentration of O2 gas? A) 3.3 × 10-11 M B) 1.7 × 10-10 M C) 6.0 × 109 M D) 3.0 × 1010 M Answer: D

OpenStudy (ashley1nonly):

for this one would we use kc= kp/ (RT)^delta n

OpenStudy (ashley1nonly):

since we are looking for the gas it would be k[=Kc(RT) delta n

OpenStudy (ashley1nonly):

Kp for the first

OpenStudy (anonymous):

kk, you would use that if given kc to find ksp,

OpenStudy (anonymous):

let me work it out real quickly, i do not want to tell you wrong

OpenStudy (ashley1nonly):

ok

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

1.) kc=(NO)^2/(N2)(O2)2 2.) 8.3*10^-10=(5)^2/(N2)(x)2 3.) 8.3*10^-10=25/N2 N2=3.0*1010M

OpenStudy (anonymous):

Does it make sense?

OpenStudy (ashley1nonly):

i get step one for part 2 i see why you plugged in kc but i dont understand everything else

OpenStudy (anonymous):

kk, We are solving it like an algebraic expression we know kc value, and it is asking what is N2(M)\ we know equilibrium is products/ reactants kc=product/reactants we were given value for NO2=5, so O2 was just 1, because it stated NO2 was 5 times of O2 so kc=8.3*10^-10, NO2=5, O2=1, N2= we are solving for 8.3*10^-10=(NO2=5)^2/(N2)(O2=1)^2 so rearrange the equation to say N2=25/8.3*10^-10= 3.0*10^-10

OpenStudy (ashley1nonly):

ok that makes a lot more sense, at first i didnt understand why o2 just left

OpenStudy (anonymous):

Aweseome, for chem 2 you can always make the other variable 1 in equilibrium

OpenStudy (ashley1nonly):

do you have time for one more

OpenStudy (anonymous):

Yeah, Im on fall break

OpenStudy (ashley1nonly):

yayyyyy lets do this okay i will start another question

OpenStudy (anonymous):

kk